OpenStudy (anonymous):

have to take a test with equations like this... 12x+4y=-1 followed by y=-3x+1 can anyone help me solve this so i can do other problems similar to it?

OpenStudy (anonymous):

What are you solving for? If you are solving for x and y, then you can't for this system of equations because they are parallel lines/have the same slope.

OpenStudy (anonymous):

yes I am solving for x and y. Thank you for your help. I didnt know they were parallel. I need to learn how to find x and y in a problem like this for my test!!

OpenStudy (anonymous):

An example problem would be 12x+4y=0 y=2x-10 substitute y from the second equation into the first equation 12x+4(2x-10)=0 solve for x 12x+8x-40=0 20x-40=0 20x=40 x=2 plug it back into the second equation y=2(2)-10 y=4-10=-6 anwer is (2,-6)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (precal):